# Revisiting Shooting Point Monte Carlo Methods for Transition Path Sampling

### v1

@@ II.5 One-way and Two-way TPS
In fixed- or flexible-length regular TPS, one draws the shooting index $k$ freshly at each round. As a result, there is no dependence of the picked candidate shooting index on the previous shooting index $k^{\text{o}}$ and Markovianity is guaranteed without the extended space formalism. Nonetheless, we can construct a move that would be non-Markovian if the shooting index was not tracked yet stays close to the original two-way shooting algorithm. For that, the algorithm can be split into two parts: one move that redraws the shooting index and the shooting move itself (Fig. 1). This way, even though the difference is subtle, a memory of the shooting index is maintained from trial to trial as the shooting move is initiated from the shooting index of the previous trial. As a result, the distribution of shooting points follows Eq. (14). For redrawing the shooting point, we replace the old shooting index $k^{\text{o}}$ with an index $k^{\text{n}}$ independently sampled from the uniform distribution:
@@
This move naturally does not require an acceptance step since it conserves the distribution of $k$ on the old path $X^{\text{o}}$. We then set the drawn shooting index to $k^{\text{o}}$ for the next move. In the second step, a shooting move is initiated that shoots from the point at $k^{\text{o}}$ and produces the path $X^{\text{n}}$. We then denote the index of the common point of $X^{\text{o}}$ and $X^{\text{n}}$ on the new path as $k^{\text{n}}$. Here, the mapping $k^{\text{o}}$ to $k^{\text{n}}$ is unique and reversible with unit generation probability. We derive the acceptance criterion based on Eq. II.3 and Eq. 17 as
@@
which is identical to the acceptance criterion derived previously Dellago et al. (2002); Juraszek and Bolhuis (2006).
In the same way, the shooting range algorithm Jung et al. (2017) can be formulated in the extended ensemble. In that case, shooting indices $k$ are only selected in a range defined by a collective variable $r(x)$:
$p(k|X)=\frac{h(r(x_{k}))}{\sum_{i}h(r(x_{i}))}\,,$ (19)
where $h(r)$ is unity if $a\leq r(x)\leq b$ with user defined bounds $a$ and $b$ and zero otherwise. The derived acceptance includes the above shooting index distribution and the condition must hold that all paths have at least one point within the shooting range Jung et al. (2017). Alternatively, a weight function $w$ can be defined for each $k$ to bias the shooting index selection
@@ II.6 Fixed-length Aimless Shooting
Aimless shooting with fixed path lengths is historically the basis for algorithms that evolve the shooting point from trial to trial. While there are different formulations of the algorithm in literature, we focus here on the two-point variant described in reference Peters (2010). Here, we start with an initial path $X^{\text{o}}$ of fixed path length $L$. The shooting index distribution is only nonzero at two-points, $L/2-\Delta k$ and $L/2+\Delta k$, where $\Delta k$ is an adjustable parameter:
@@
The two-point aimless algorithm starts by selecting one of these two indices on the old path as a shooting index, which one can describe as an index redrawing as above for regular TPS. Then, integration is performed in the backward direction by $L/2-\Delta k$ and in the forward direction by $L/2+\Delta k$ to obtain $X^{\text{n}}$. This produces a new path on which the index of the common point $\hat{k}^{\text{n}}$ can be either at index $L/2-\Delta k$ if the path goes from state A to state B or at index $L/2+\Delta k$ for a state B to state A path. In equilibrium, both of these outcomes have equal probability and therefore the new shooting index follows the desired $p(k|X)$. Due to the splitting of the generation of the new shooting index and the shooting move itself, the acceptance for the shooting move only includes the shooting index distributions
@@
$=H_{\text{AB}}(X^{\text{n}})\min\biggl{[}1,\frac{p(\hat{k}^{\text{n}}|X^{\text{n}})}{p(\hat{k}^{\text{o}}|X^{\text{o}})}\biggr{]}$
@@
It is important to note that the shooting point does not evolve in fixed-length aimless shooting compared to its flexible-length counterpart. While there is a chance that the next shooting move is initiated from the same point as in the last trial, the shooting index is always drawn independently as $L/2-\Delta k$ or $L/2+\Delta k$ each trial, effectively erasing the memory of the previous shooting index. The distribution of the shooting points from which a shoot is attempted is according to Eq. II.4:
$+\delta(x_{L/2+\Delta k}-x)]\,,$ (24)
which is the distribution of points ±$\pm\Delta k$ away from the path centers. We validate this result via numerical simulations in section III.
@@ II.7 Flexible-length Aimless Shooting
Flexible-length aimless shooting, as proposed by Mullen et al. Mullen et al. (2015), generalizes the aimless shooting move to a flexible-length scenario, which is particularly useful in systems with diffusive dynamics. Similar to fixed-length aimless shooting, the shooting index for the following trial is chosen from a narrow set of possible indices. Due to the flexible-length setting, these points are no longer defined around the path center as in Eq. 22, but close to the previous successful shooting point. The proposed acceptance criterion does not include a ratio of path lengths Mullen et al. (2015).In the following, we will show, based on our new formalism, that such a factor is necessary to sample the correct transition path ensemble.
Using the proposed shooting point Monte Carlo formalism, we see a crucial change going from fixed- to flexible-length moves. Namely, the new shooting index $k^{\text{n}}$ is selected depending on the previous index $k^{\text{o}}$ (Fig. 2). In particular, the two-point variant described in Mullen et al. (2015) samples the new shooting index from $\{k^{\text{o}},k^{\text{o}}+s\Delta k\}$, where $s\in\{-1,1\}$ defines whether the second shooting point is on the backward or forward segment. Since the sign $s$ is preserved until a next trial is accepted, we include it in the extended ensemble $(X,k,s)$ (see SI for detailed balance derivation). To respect reversibility, the generation probability of $k^{\text{n}}$ is divided into two steps. We first choose a shooting index $\hat{k}^{\text{o}}$ on the old path according to:
@@ II.8 Spring Shooting
Inspired by the flexible-length aimless shooting scheme, the work by Brotzakis et al. Brotzakis and Bolhuis (2016) aimed to achieve a similar efficiency gain in a one-way shooting framework. The displacement $\Delta k$ from the previous shooting index is chosen according to an exponential distribution, i.e., a Boltzmann distribution based on a linear potential (constant force), hence the scheme is named spring shooting. We reformulate the algorithm in the shooting point MC framework (Fig. 3), starting out by selecting the direction of the shooting move and setting $s=-1$ for a forward and $s=1$ for a backward shoot. Note that in this case, $s$ is redrawn at the start of every trial and therefore does not have to be included in the extended space. The displacement $\Delta k$ is then sampled according to:
@@
where $\sigma$ is the strength of the spring and $\Delta k_{\text{max}}$ is the maximum displacement. Following the notation of flexible-length aimless shooting, the shooting index $\hat{k}^{\text{o}}$ is set as $k^{\text{o}}+\Delta k$. The generation probability of the new shooting index becomes
@@
Then, the new path $X^{\text{n}}$ is generated from the configuration at $\hat{k}^{\text{o}}$. The index of the common point of $X^{\text{n}}$ and $X^{\text{o}}$ on the new path is denoted as $\hat{k}^{\text{n}}$. As for aimless shooting, we then shift the new shooting index $\hat{k}^{\text{n}}$ a second time by sampling a shift $\Delta k$ from
@@
This means that, given a forward shot, first the shooting index is shifted preferably backwards and, after generating the new path, the new shooting index is generated by shifting preferably forward. Thereby, the move is reversible and generation probabilities are symmetric (see SI). In this case, the resulting acceptance ratio is identical to Eq. 28 and a shooting index distribution must be defined.
Figure 3: The spring shooting move in the extended space of paths $X$ and shooting indices $k$. (1) The move starts with the current path $X^{\text{o}}$ and shooting index $k^{\text{o}}$. (2) The first index shift generating $\hat{k}^{\text{o}}$ is performed by a weighted selection from neighboring points (illustrated using color gradient). (3) An extended space shooting move in the forward direction generates $X^{\text{n}}$ and $\hat{k}^{\text{n}}$. (4) A second index shift is performed leading to $k^{\text{n}}$. This shift is performed with opposite weights (see Eq. 31) and is required to ensure the reverse move is possible (tracing bottom to top).
In the original algorithm proposed by Brotzakis et al. Brotzakis and Bolhuis (2016), the new shooting index $k^{\text{n}}$ is set implicitly to $\hat{k}^{\text{n}}$, the common point of the new an old path, i.e., the second shifting is omitted. While this formulation of spring shooting allows the reverse move in path space, i.e., regenerating the old path $X^{\text{o}}$ from the new path $X^{\text{n}}$, it does not regenerate the old shooting index $k^{\text{o}}$ from $k^{\text{n}}$. As a result, detailed balance is violated in the extended space framework. As for flexible-length aimless shooting, the ratio of shooting index distributions is not considered in the original form of the sampling scheme. In the case of spring shooting, a reweighting factor for paths sampled without a defined shooting index distribution cannot be derived due to the irreversibility of the shooting move in the extended space.
@@ III Results and Discussion
Figure 5: Flexible-length transition path sampling using aimless shooting as proposed by Mullen et al. Mullen et al. (2015) and spring shooting as proposed by Brotzakis et al. Brotzakis and Bolhuis (2016) (top row) and with the proposed corrections (bottom row). (A, D) Comparison of the density of points on paths between the equilibrium reference (dashed line) and path sampling algorithms. (B,E) Distribution of points at $k$ in the $(X,k)$ extended ensemble for the different path sampling algorithms. The reference distribution was obtained via reweighting the reference density in panel A with a factor proportional to $1/L(X)$. (C, F) The distribution of path lengths obtained using the different path sampling schemes.
@@
For the fixed-length path ensemble (Fig. 4B,C), we compare the density of points on paths $\rho(x|\text{TP})$ as in Eq. II.4 with the shooting point distribution as defined in Eq. II.4. The density of points in the transition region between states A and B decreases for longer paths as less time is spent in the barrier region. This in turn leads to the points at $L/2\pm\Delta k$ being less constrained to be on the barrier top as the time at which the path leaves state A becomes increasingly variable (Fig. 4C). For the longest paths in these simulations, there is hardly any difference between the density of points at $L/2\pm\Delta k$ and $\rho(x|\text{TP})$. Since aimless shooting in a fixed-length setting samples these points as shooting points, the performance of the shooting scheme will depend on the chosen path length. For long paths, both uniform, two-way and aimless shooting initiate shooting moves effectively from the same shooting point distribution. As the shooting index is independently drawn at each trial and the shooting point distribution follows Eq. II.6, the efficiency of fixed-length aimless shooting originates from shooting close to path centers rather than from a restoring force on the shooting point that directs it to the barrier top. For long path lengths, the acceptance of the shooting move suffers as shooting points are primarily taken from the stable states rather than the barrier region.
@@
Continuing with the flexible-length path ensemble, we see that uniform two-way shooting converges to the reference ensemble, indicated by overlapping $\rho(x|\text{TP})$ distributions and path length distributions $p(L)$ (Fig. 5). The density of points at $k$ noted as $\rho_{\text{ens}}(x)$ matches the reweighted density of points on paths (Eq. 15) obtained from the reference trajectory since shooting indices are selected with a uniform $p(k|X)$ on each path. In contrast, flexible-length aimless shooting as proposed by Mullen et al. Mullen et al. (2015) and spring shooting as proposed by Brotzakis et al. Brotzakis and Bolhuis (2016) do not converge to the reference, as can be seen in panels A and C of Fig. 5. Panel B shows that the shooting points for spring shooting are much more focused around the location of the barrier. While this is a desired feature, it clearly gives rise to the wrong distribution of transition paths (panels A and C). We note that the deviations from the reference path ensemble are less clearly seen at higher barriers, where the distribution of shooting points is more localized for all methods. This is also the reason why the spring shooting led to seemingly converged results in Brotzakis and Bolhuis (2016).
@@
Including the shooting index distribution $p(k|X)$ in the acceptance and, for spring shooting, performing a second shift of the shooting index, recovers the convergence to the reference path ensemble (panels D, E, and F of Fig. 5). However, since we impose a uniform $p(k|X)$, shooting moves are now initiated from the same points as in standard uniform two-way shooting. In addition, the small displacements of the shooting index in aimless and spring shooting introduce correlations between shooting points that are absent in uniform one- or two-way shooting. These correlations lead to a slow exploration of shooting points away from the barrier top. For sufficiently large barriers, this can lead to disproportionate shooting from the barrier top in the early stages of sampling, which in turn leads to a slow convergence to the correct path ensemble.

### v2

@@ II.5 One-way and Two-way TPS
We will next cast various TPS algorithms in the extended space formalism, starting with regular one-way and two-way shooting. In fixed- or flexible-length regular TPS, one draws the shooting index $k$ freshly at each round. As a result, there is no dependence of the picked shooting index on the previous shooting index $k^{\text{o}}$ and Markovianity is guaranteed without the extended space formalism. In fact, since regular TPS is carried out in trajectory space only, the acceptance rules derived for the extended space do not directly apply to this case. Vice versa, casting the algorithm unmodified to the extended space leads to irreversibility of the move (Fig. 1). However, we can construct an alternative TPS scheme that (a) is Markovian in the extended space, and (b) closely resembles the original two-way shooting algorithm.
We split the shooting move into three substeps as illustrated in Fig. 2. In the first part, the index shift, the shooting index is redrawn shifting the old shooting point $k^{\text{o}}$ to the point $\hat{k}^{\text{o}}$. In the second step, the shooting move itself is carried out from the point at $\hat{k}^{\text{o}}$. As the shooting move in most cases leads to a renumbering of points, we then define $\hat{k}^{\text{n}}$ as the index of the shooting point on the new path. As a last step, the shooting index is redrawn once again to ensure reversibility of the move, resulting in a new path and shooting index $(X^{\text{n}},k^{\text{n}})$.
Since we now work in the extended space, we need to specify the target shooting index probability $p(k|X)$ in order to obtain the full joint probability $P(X,k)$. To enforce a uniform shooting index distribution, we define $p(k|X)$ as:
@@
For both index shift moves (Fig. 2), we draw the shooting index from the uniform distribution to simplify the acceptance and keep the algorithm close to standard TPS, such that
$p_{\text{sel}}(\hat{k}^{\text{o}}|X^{\text{o}})=\frac{1}{L(X^{\text{o}})}\ \text{for}\ 1\leq k\leq L(X^{\text{o}})$
$p_{\text{sel}}(\hat{k}^{\text{n}}|X^{\text{n}},X^{\text{o}},\hat{k}^{\text{o}})=\frac{1}{L(X^{\text{n}})}\ \text{for}\ 1\leq k\leq L(X^{\text{n}})\,.$ (18)
Combining both shifts leads to a symmetric generation probability for the forward and the reverse move:
$p_{\text{gen}}(k^{\text{n}}|X^{\text{n}},X^{\text{o}},k^{\text{o}})$ $=p_{\text{gen}}(k^{\text{o}}|X^{\text{o}},X^{\text{n}},k^{\text{n}})$
$=\frac{1}{L(X^{\text{n}})}\frac{1}{L(X^{\text{o}})}\,.$ (19)
After cancellation of the shooting index generation probabilities defined above, the acceptance probability based on Eqs. (II.3) and (17) is then given by
@@
which is identical to the acceptance criterion derived previously Dellago et al. (2002); Juraszek and Bolhuis (2006). Note that for this algorithm the distribution of shooting points $\rho(x|\text{SP})$ follows Eq. (15).
Figure 2: General scheme of a reversible shooting move in the extended space. The shooting index is shifted and the equations of motion are integrated. After that, the new shooting index is shifted a second time ensure reversibility.
In the analogous way, the shooting range algorithm Jung et al. (2017) can be formulated in the extended ensemble. In this case, shooting indices $k$ are only accepted in a range defined by a collective variable $r(x)$, corresponding to a shooting index distribution
$p(k|X)=\frac{h(r(x_{k}))}{\sum_{i}h(r(x_{i}))}\,.$ (21)
Here, $h(r)$ is unity if $a\leq r(x)\leq b$ with user defined bounds $a$ and $b$ and zero otherwise. The derived acceptance includes the above shooting index distribution and the condition must hold that all paths have at least one point within the shooting range Jung et al. (2017). Alternatively, a weight function $w$ can be defined for each $k$ to bias the shooting index selection
@@ II.6 Fixed-length Aimless Shooting
Aimless shooting with fixed path lengths is historically the basis for algorithms that evolve the shooting point from trial to trial. While there are different formulations of the algorithm in literature, we focus here on the two-point variant described in Ref. Peters (2010). Here, we start with an initial path $X^{\text{o}}$ of fixed path length $L$. The shooting index distribution is only nonzero at two points, $L/2-\Delta k$ and $L/2+\Delta k$, where $\Delta k$ is an adjustable parameter. This choice corresponds to the target shooting index distribution
@@
The two-point aimless shooting algorithm starts by selecting one of these two indices on the old path as the shooting index $\hat{k}^{\text{o}}$, which corresponds to the first index shift for regular TPS (Figure 2). Then, integration is performed in the backward direction by $L/2-\Delta k$ steps and in the forward direction by $L/2+\Delta k$ steps to obtain the new path $X^{\text{n}}$. This operation produces a new path on which the index of the common point, i.e., the index $\hat{k}^{\text{n}}$ of the point where the old and the new path intersect, can be either at $L/2-\Delta k$ or at $L/2+\Delta k$.
In principle, a second shift from $\hat{k}^{\text{n}}$ to $k^{\text{n}}$ would now be required to ensure that the move is reversible in the extended space. However, it is important to note that in fixed-length aimless shooting the shooting point does not evolve, in contrast to to its flexible-length counterpart. While there is a chance that the next shooting move is initiated from the same point as in the last trial, the shooting index is always drawn independently as $L/2-\Delta k$ or $L/2+\Delta k$ each trial, effectively erasing the memory of the previous shooting index. In other word, given a certain path $X^{\text{o}}$, determining the new shooting index is always done relative to the center $L/2$ of the path and does not require knowledge of the previous shooting index. As a result, the procedure is Markovian in path space and therefore does not require an extended space formalism with second shift. We introduce a second shift as for standard TPS, which allows us to apply the extended space formalism and infer the distribution of the shooting points from which a shooting is attempted via Eq. II.4:
$+\delta(x_{L/2+\Delta k}-x)]\,.$ (25)
This is the distribution of points ±$\pm\Delta k$ away from the path centers. We validate this result via numerical simulations in section III. As for standard TPS in extended space, performing both shifts with symmetric probabilities leads to a cancellation in the acceptance probability
@@ II.7 Flexible-length Aimless Shooting
Flexible-length aimless shooting, as proposed by Mullen et al. Mullen et al. (2015), generalizes the aimless shooting move to a flexible-length scenario, which is particularly useful in systems with diffusive dynamics. Similar to fixed-length aimless shooting, the shooting index for the following trial is chosen from a narrow set of possible indices. Due to the flexible-length setting, these points are no longer defined to lie around the path center as in Eq. 24, but they are close to the previous shooting point. Hence, in aimless shooting with flexible length pathways, there is true memory of the previous shooting point, making the treatment of the algorithm in the extended space formalism necessary. In the following, we will formulate the flexible-length aimless shooting in this formalism and will show that the acceptance probability include a ratio of path lengths, which was neglected in the original work Mullen et al. (2015). As demonstrated below numerically, this factor is necessary to sample the correct transition path ensemble.
As mentioned above, there is a crucial difference between the fixed-length and the flexible-length moves. Namely, in flexible-length algorithm, the new shooting index $k^{\text{n}}$ is selected depending on the previous index $k^{\text{o}}$ (see Fig. 3). In particular, the two-point variant described in Mullen et al. (2015) samples the new shooting index from $\{k^{\text{o}},k^{\text{o}}+s\Delta k\}$, where $s\in\{-1,1\}$ defines whether the second shooting point is on the backward or the forward trajectory segment. Since the sign $s$ is preserved until a next trial is accepted, we include it in the extended ensemble $(X,k,s)$ (see SI for detailed balance derivation). To respect reversibility, the generation probability of $k^{\text{n}}$ is divided into two steps. We first choose a shooting index $\hat{k}^{\text{o}}$ on the old path according to
@@ II.8 Spring Shooting
Inspired by the flexible-length aimless shooting scheme, the spring shooting algorithm of Brotzakis et al. Brotzakis and Bolhuis (2016) aimed to achieve a similar efficiency gain in a one-way shooting framework. In this scheme, the displacement $\Delta k$ from the previous shooting index is chosen according to an exponential distribution, i.e., a Boltzmann distribution based on a linear potential (constant force), hence the scheme is named spring shooting. We next reformulate the algorithm in the shooting point MC framework (Fig. 4), beginning by selecting the direction of the shooting move and assigning $s=-1$ for a forward shooting and $s=1$ for a backward shooting. Note that in this case $s$ is drawn freshly at the start of every trial, so it does not need to be included in the extended space. The displacement $\Delta k$ is then sampled according to:
@@
where $\sigma$ is the strength of the spring and $\Delta k_{\text{max}}$ is the maximum displacement. Following the notation of flexible-length aimless shooting, the shooting index $\hat{k}^{\text{o}}$ is set to $k^{\text{o}}+\Delta k$. The generation probability of the new shooting index becomes
@@
Then, the new path $X^{\text{n}}$ is generated starting from the configuration at $\hat{k}^{\text{o}}$. The index of the common point of $X^{\text{n}}$ and $X^{\text{o}}$ on the new path is denoted as $\hat{k}^{\text{n}}$. As for aimless shooting, we then shift the new shooting index $\hat{k}^{\text{n}}$ a second time by sampling a shift $\Delta k$ from
@@
This implies that, for a forward shot, first the shooting index is shifted preferably backwards. After generating the new path, the new shooting index is then generated by shifting, preferably forward. This ensures the move is reversible, and generation probabilities are symmetric (see SI). Consequently, the resulting acceptance ratio is identical to Eq. 30 and a distribution for the shooting index must be defined.
Figure 4: The spring shooting move in the extended space of paths $X$ and shooting indices $k$. (1) The move starts with the current path $X^{\text{o}}$ and shooting index $k^{\text{o}}$. (2) The first index shift generating $\hat{k}^{\text{o}}$ is performed by a weighted selection from neighboring points (illustrated using color gradient). (3) An extended space shooting move in the forward direction generates $X^{\text{n}}$ and $\hat{k}^{\text{n}}$. (4) A second index shift is performed leading to $k^{\text{n}}$. This shift is performed with opposite weights (see Eq. 33) and is required to ensure that the reverse move is possible (tracing bottom to top).
In the original spring shooting algorithm proposed by Brotzakis et al. Brotzakis and Bolhuis (2016), the new shooting index $k^{\text{n}}$ is set implicitly to $\hat{k}^{\text{n}}$, the common point of the new and old path, i.e., the second shifting is omitted. While this formulation of spring shooting allows the reverse move in path space, i.e., regenerating the old path $X^{\text{o}}$ from the new path $X^{\text{n}}$, it does not regenerate the old shooting index $k^{\text{o}}$ from $k^{\text{n}}$. As a result, detailed balance is violated in the extended space framework. As for flexible-length aimless shooting, the ratio of shooting index distributions is not considered in the original form of the sampling scheme. For spring shooting, however, a reweighting factor for paths sampled without a defined shooting index distribution cannot be derived, due to the irreversibility of the shooting move in the extended space.
Figure 6: Flexible-length transition path sampling using aimless shooting as proposed by Mullen et al. Mullen et al. (2015) and spring shooting as proposed by Brotzakis et al. Brotzakis and Bolhuis (2016) (top row) and with the proposed corrections (bottom row). (A, D) Comparison of the density of points on paths between the equilibrium reference (dashed line) and path sampling algorithms. (B,E) Distribution of points after integrating out $k$ in the $(X,k)$ extended ensemble for the different path sampling algorithms. The reference distribution was obtained via reweighting the reference density in panel A with a factor proportional to $1/L(X)$. (C, F) The distribution of path lengths obtained using the different path sampling schemes.
@@ III Results and Discussion
For the fixed-length path ensemble (Fig. 5B,C), we compare the density of points on paths $\rho(x|\text{TP})$ as in Eq. II.4 with the shooting point distribution as defined in Eq. II.4. The density of points in the transition region between states A and B decreases for longer paths as less time is spent in the barrier region. As a result, the points at $L/2\pm\Delta k$ are less constrained to remain at the barrier top, as the time at which the path leaves state A becomes increasingly variable (Fig. 5C). For the longest paths in these simulations, there is hardly any difference between the density of points at $L/2\pm\Delta k$ and $\rho(x|\text{TP})$. Since aimless shooting in a fixed-length setting samples these points as shooting points, the performance of the shooting scheme will depend on the chosen path length. For long paths, both uniform, two-way and aimless shooting initiate shooting moves effectively from the same shooting point distribution. As the shooting index is drawn independently at each trial and the shooting point distribution follows Eq. II.6, the efficiency of fixed-length aimless shooting arises from shooting close to path centers, rather than from any restoring force that directs shooting point toward the barrier top. For long path lengths, the acceptance probability of the shooting move decreases, as shooting points are more frequently drawn from the stable states rather than the barrier region.
@@
Within the flexible-length path ensemble, we observe that uniform two-way shooting converges to the reference ensemble, as indicated by overlapping $\rho(x|\text{TP})$ distributions and path length distributions $p(L)$ (Fig. 6). The density of points at $k$ denoted as $\rho_{\text{ens}}(x)$ matches the reweighted density of points on paths (Eq. 15) obtained from the reference trajectory since shooting indices are selected with a uniform $p(k|X)$ on each path. In contrast, flexible-length aimless shooting as proposed by Mullen et al. Mullen et al. (2015) and spring shooting as proposed by Brotzakis et al. Brotzakis and Bolhuis (2016) do not converge to the reference, as shown in panels A and C of Fig. 6. Panel B shows that the shooting points for spring shooting are much more focused around the location of the barrier. While this is a desired feature, it clearly gives rise to the wrong distribution of transition paths (panels A and C). We note that the deviations from the reference path ensemble are less apparent at higher barriers, where the distribution of shooting points is more localized for all methods. This is also the reason why spring shooting appeared to produce seemingly converged results in Brotzakis and Bolhuis (2016).
@@
Including the shooting index distribution $p(k|X)$ in the acceptance probability and, for spring shooting, performing a second shift of the shooting index, recovers the convergence to the reference path ensemble (panels D, E, and F of Fig. 6). However, since we impose a uniform $p(k|X)$, shooting moves are now initiated from the same points as in standard uniform two-way shooting. In addition, the small displacements of the shooting index in aimless and spring shooting introduce correlations between shooting points that are absent in uniform one- or two-way shooting. These correlations result in slow exploration of shooting points away from the barrier top. For sufficiently large barriers, this can cause disproportionate shooting from the barrier top during the early stages of sampling, which in turn leads to a slow convergence to the correct path ensemble.